# Counting integer matrices with square-free determinants

### v1

@@ Abstract.
We consider the set ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ of $n\times n$-matrices with integer elements of size at most $H$ and obtain and asymptotic formula on the number of matrices from ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ with square-free determinants. We also use our approach with some further enhancements, to obtain an asymptotic formula for the sum of the Euler function with determinants of matrices from ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$.
@@ 1. Introduction
1.1. Motivation and main results
For a positive integer $n$, we use ${\mathcal{M}}_{n}\left({\mathbb{Z}}\right)$ to denote the set of all $n\times n$ matrices with integer elements. Furthermore, for a real $H\geq 1$ we use ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ to denote the set of matrices
@@ 1.1. Motivation and main results
with integer entries of size $|a_{ij}|\leq H$. In particular, ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ is of cardinality #$\#{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)=\left(2\left\lfloor H\right\rfloor+1\right)^{n^{2}}$.
Motivated by recent investigation of some arithmetic properties of matrix determinants and minors, see [9, 10], here we obtain an asymptotic formula for the number $S_{n}(H)$ of matrices $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ with a square-free determinant, that is, matrices for which $\det A$ is not divisible by a square of any prime.
We remark that treating the determinant as a generic polynomial over ${\mathbb{Z}}$ in $n^{2}$ variables and of degree $n$, is not going to bring any results as at the present time all results on square-free values of polynomials of high degree are conditional on the celebrated $abc$-conjecture, see [11, 12].
Our asymptotic formula gives a power saving in the error term, exceeding $1/2$. In fact, using the standard inclusion-exclusion principle together with Lemmas 2.1 and 2.2 below one can already get a power saving in the error term. However we go beyond this and get a stronger bound.
@@
$S_{n}(H)=\mathfrak{S}_{n}H^{n^{2}}+O\left(H^{n^{2}-\gamma_{n}}\right),$
@@
$\mathfrak{S}_{n}=\prod_{p\leavevmode\nobreak\ \text{prime}}\,\prod_{j=2}^{n+1}\left(1-p^{-j}\right)$
@@
Our approach is based on new bounds on exponential sums along determinant hypersurfaces, which we believe could be of independent interest. We emphasise that the main feature of Theorem 1.1 is that $\gamma_{n}>1/2$ since with $\gamma_{n}=n^{2}/(2n^{2}+2)<1/2$ it can be derived via the aforementioned elementary argument, avoiding the use of exponential sums.
@@
$\varPhi_{n}(H)=\sum_{\begin{subarray}{c}A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)\\ \det A\neq 0\end{subarray}}\frac{\varphi\left(|\det A|\right)}{|\det A|}.$
@@
$\varPhi_{n}(H)=\sigma_{n}H^{n^{2}}+O\left(H^{n^{2}-\vartheta_{n}+o(1)}\right),$
@@
$\sigma_{n}=\prod_{p\leavevmode\nobreak\ \text{prime}}\,\left(1-\frac{1}{p}\right)\left(1+\frac{1}{p}\prod_{j=2}^{n}\left(1-p^{-j}\right)\right)$
@@
Our proof of Theorem 1.2 takes advantage of stronger bounds of exponential sums with square-free denominators with so-called monomial linear forms, see Lemma 3.7 below. We have not been able to establish an analogues of this bound in the case of denominators which are squares of square-free integers.
@@ 1.2. Notation and conventions
For a finite set ${\mathcal{S}}$ we use #$\#{\mathcal{S}}$ to denote its cardinality.
@@
as $|k|\to\infty$, see [6, Equation (1.81)].
We also use the convention that if $A$ is an $n\times n$ matrix defined over ${\mathbb{Z}}_{m}$ or ${\mathbb{F}}_{p}$, then by $\det A=0$ we mean $\det A\equiv 0\bmod m$ or $\det A\equiv 0\bmod p$, respectively, which should be clear from the context. When $A$ has entries in an interval $[-H,H]$, then any condition on the determinant modulo a positive integer $m$ or prime $p$ will be clearly indicated.
@@ 2.2. Matrices with fixed determinants
We need a bound on the number of matrices $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ with prescribed value of the determinant $\det A=d$. We recall that Duke, Rudnick and Sarnak [3], if $d\neq 0$, and Katznelson [7], when $d=0$, have obtained asymptotic formulas (with the main terms of orders $H^{n^{2}-n}$ and $H^{n^{2}-n}\log H$, respectively) for the number of such matrices when $d$ is fixed. However this is too restrictive for our purpose and so we use a uniform with respect to $d$ upper bound which is a special case of [13, Theorem 4].
@@ 3.1. Set-up
We are only interested in the sums $S_{m}(L)$ when the modulus $m=d$ or $m=d^{2}$ for a square-free integer $d\geq 1$. Thus we start with reduction to sums modulo $p$ or $p^{2}$. Namely, by the Chinese Remainder Theorem, we have the following decomposition, see [6, Equation (12.21)].
@@ 3.2. Linear sections of the variety of singular matrices
[…]
@@
[…]
@@ 3.3. Bounding $S_{p}(L)$
Combining Lemma 3.2 with bounds of exponential sums along algebraic varieties, see, for example, [4, Proposition 1.2], one can immediately derive
@@
[…]
@@
for some linear forms $\ell$ and $L^{*}$ in $n$ and $n(n-1)$ variables, respectively, where, by our assumption, $\ell$ is nontrivial modulo $p$.
@@ 3.4. Bounding $S_{p^{2}}(L)$
for some linear forms $\ell$ and $L^{*}$ in $n$ and $n(n-1)$ variables, respectively, where by our assumption, $\ell$ is nontrivial modulo $p$.
@@
[…]
@@ 3.5. Bounding $S_{d}(L)$ and $S_{d^{2}}(L)$
[…]
@@
[…]
@@ 4.1. Discrepancy and the Koksma-Szüsz inequality
We recall that the discrepancy $D(\Gamma)$ of a sequence $\Gamma=(\bm{\gamma}_{k})_{k=1}^{K}$ of $K$ points in the half-open $s$-dimensional unit cube $[0,1)^{s}$, is defined by
@@
One of the basic tools in the uniformity of distribution theory is the celebrated Koksma–Szüsz inequality [8, 14] (see also [2, Theorem 1.21]), which links the discrepancy of a sequence of points to certain exponential sums.
@@ 4.2. Matrices with determinants with a divisibility condition in boxes
$N_{n}(m;H)=\#\{A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right):\leavevmode\nobreak\ \det A\equiv 0\bmod m\}.$
@@
$N_{n}(d;H)=N_{n}(d)\frac{H^{n^{2}}}{d^{n^{2}}}+O\left(d^{n^{2}-2+1/n+o(1)}\right).$
@@
Let $\mathbf{H}=\left(H_{ij}\right)_{i,j=1}^{n}$ with some positive integers $H_{ij}\leq m$. We extend the definition of $N_{n}(m;H)$ to $N_{n}(m;\mathbf{H})$ which is the number of $n\times n$ matrices with
$A=\left(a_{ij}\right)_{i,j=1}^{n}\in{\mathcal{M}}_{n}\left({\mathbb{Z}}\right),\qquad 0\leq a_{ij}<H_{ij},\ i,j=1,\dots,n,$
@@
(4.1) $N_{n}(d;\mathbf{H})=N_{n}(d)\frac{1}{d^{n^{2}}}\prod_{i,j=1}^{n}H_{ij}+O\left(d^{n^{2}-2+1/n+o(1)}\right).$
@@
For a positive integer $M\geq 1$, it is convenient to define ${\mathcal{L}}(M)$ as the set of all nonzero linear forms $L(\mathbf{X})\in{\mathbb{Z}}[\mathbf{X}]$ of the form (3.1) and with coefficients in the interval $[-M,M]$.
@@
Hence, splitting each interval $[0,H_{ij})$ into intervals of length $d$, and applying Lemma 4.1 to count points in incomplete boxes on the boundary, we have
$N_{n}(d;\mathbf{H})=N_{n}(d)\frac{1}{d^{n^{2}}}\prod_{i,j=1}^{n}H_{ij}+O\left(M^{-1}N_{n}(d)+R\right),$
@@
(4.2) $N_{n}(d)\leq d^{n^{2}}\prod_{p\mid d}\left(p^{-1}+O\left(p^{-2}\right)\right){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}=}d^{n^{2}-1+o(1)}.$
@@
(4.3) $N_{n}(d;\mathbf{H})=N_{n}(d)\frac{1}{d^{n^{2}}}\prod_{i,j=1}^{n}H_{ij}+O\left(d^{n^{2}-1+o(1)}M^{-1}+R\right).$
@@
$\sum_{\begin{subarray}{c}L\in{\mathcal{L}}_{d}(M)\\ L\leavevmode\nobreak\ \text{monomial}\\ \gcd(L,d)=D\end{subarray}}\frac{1}{\rho(L)}\ll D^{-1}\log d.$
@@
(4.5) $R_{\text{mon}}=d^{n^{2}-n+o(1)}\sum_{\begin{subarray}{c}D\mid d\\ D\leq M\end{subarray}}D^{n}\sum_{\begin{subarray}{c}L\in{\mathcal{L}}_{d}(M)\\ L\leavevmode\nobreak\ \text{monomial}\\ \gcd(L,d)=D\end{subarray}}\frac{1}{\rho(L)}\leq d^{n^{2}-n+o(1)}M^{n-1}.$
@@
$\sum_{\begin{subarray}{c}L\in{\mathcal{L}}_{d}(M)\\ L\leavevmode\nobreak\ \text{non-monomial}\\ \gcd(L,d)=D\end{subarray}}\frac{1}{\rho(L)}\ll D^{-2}(\log d)^{n^{2}}.$
@@
(4.6) $\begin{split}R_{\text{non-mon}}&=d^{n^{2}-(n+1)/2+o(1)}\sum_{\begin{subarray}{c}D\mid d\\ D\leq M\end{subarray}}D^{(n+1)/2}\sum_{\begin{subarray}{c}L\in{\mathcal{L}}_{d}(M)\\ L\leavevmode\nobreak\ \text{non-monomial}\\ \gcd(L,d)=D\end{subarray}}\frac{1}{\rho(L)}\\ &\leq d^{n^{2}-(n+1)/2+o(1)}\left(M^{(n-3)/2}+1\right).\end{split}$
@@
$N_{n}(d;\mathbf{H})=N_{n}(d)\frac{1}{d^{n^{2}}}\prod_{i,j=1}^{n}H_{ij}+O\left(d^{n^{2}+o(1)}\mathscr{E}\right),$
@@
[…]
@@
$N_{n}(d^{2};H)=N_{n}(d^{2})\frac{H^{n^{2}}}{d^{2n^{2}}}+O\left(d^{2n^{2}-4(n+1)/(n+3)+o(1)}\right).$
@@
We follow the same path as in the proof of Lemma 4.2. In particular, we continue to use the notations $\mathbf{H}=\left(H_{ij}\right)_{i,j=1}^{n}$ and $N_{n}(m;\mathbf{H})$ and observe that it is enough to show that
(4.7) $N_{n}(d^{2};\mathbf{H})=N_{n}(d^{2})\frac{1}{d^{2n^{2}}}\prod_{i,j=1}^{n}H_{ij}+O\left(d^{2n^{2}-4(n+1)/(n+3)+o(1)}\right).$
@@
$N_{n}(d^{2};\mathbf{H})=N_{n}(d^{2})\frac{1}{d^{2n^{2}}}\prod_{i,j=1}^{n}H_{ij}+O\left(M^{-1}N_{n}(d^{2})+R\right),$
@@
(4.9) $N_{n}(d^{2};\mathbf{H})=N_{n}(d^{2})\frac{1}{d^{2n^{2}}}\prod_{i,j=1}^{n}H_{ij}+O\left(M^{-1}d^{2n^{2}-2+o(1)}+R\right).$
@@
$N_{n}(d^{2};\mathbf{H})$ $=N_{n}(d^{2})\frac{1}{d^{2n^{2}}}\prod_{i,j=1}^{n}H_{ij}$
@@
[…]

### v2

@@ Abstract.
We consider the set ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ of $n\times n$-matrices with integer elements of size at most $H$ and obtain an asymptotic formula for the number of matrices from ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ with square-free determinants. We also use our approach with some further enhancements, to obtain an asymptotic formula for the sum of the Euler function with determinants of matrices from ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$.
@@ 1. Introduction
1.1. Motivation results
For a positive integer $n$, we use ${\mathcal{M}}_{n}\left({\mathbb{Z}}\right)$ to denote the set of all $n\times n$ matrices with integer elements. Furthermore, for an integer $H\geq 1$ we use ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ to denote the set of matrices
@@ 1.1. Motivation results
with integer entries of size $|a_{ij}|\leq H$. In particular, ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ is of cardinality ${\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)=\left(2H+1\right)^{n^{2}}$.
Recently there has been quite active interest in investigating arithmetic properties of matrix determinants and minors, , see [11, 12]. For example, Kotsovolis and Woo [11] obtained an asymptotic formula for the number of matrices (counted by an slightly different norm) with prime determinants. The methods of homogeneous dynamics, underlying the approach of [11], do not allow to get strong bounds on the error terms of this asymptotic formula. Motivated by this work, and also by work of Kovaleva [12] on square-free minors, we obtain an asymptotic formula for the number $S_{n}(H)$ of matrices $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ with a square-free determinant, that is, matrices for which $\det A$ is not divisible by a square of any prime. While investigating square-free values is typically easier than investigating prime values, our goal is to get strong bounds on the error term of our asymptotic formula.
We also remark that treating the determinant as a generic polynomial over $\mathbb{Z}$} in $n^{2}$ variables and of degree $n$, is not going to bring any results as at the present time all results on square-free values of polynomials of high degree are conditional on the celebrated abcabc-conjecture, see [13, 14], or require an exponentially large (compared to the degree) number of variables [2].
1.2. Main results
Our asymptotic formula gives a power saving in the error term, exceeding $1/2$, which is the limit of approaches based on the use of the Weil bound.
In fact, using the standard inclusion-exclusion principle together with Lemmas 2.1 and 2.2 below one can already get a power saving in the error term. However we go beyond this and get a stronger bound.
@@ 1.2. Main results
$S_{n}(H)=2^{n^{2}}\mathfrak{S}_{n}H^{n^{2}}+O\left(H^{n^{2}-\gamma_{n}}\right),$
@@
$\mathfrak{S}_{n}=\prod_{p\penalty 10000\ \text{prime}}\,\prod_{j=2}^{n+1}\left(1-p^{-j}\right)$
@@
Our approach is based on new bounds on exponential sums along determinant hypersurfaces, which we believe could be of independent interest.
Remark 1.2.
Note that the bounds based on algebraic geometry, such as in [5, 9], do not apply to rational exponential sum modulo $p^{2}$, where $p$ is prime, which is our principal case. Moreover, even in the case of exponential sums modulo $p$, our bounds are stronger that those which can be derived from, for example, results of Katz [8]. This is due to the high dimension of the singularity locus of the determinant variety over ${\mathbb{F}}_{p}$ (which is contained in the set of matrices of $n\times n$ matrices over ${\mathbb{F}}_{p}$ of rank at most n2n-2), affecting the strength of the algebraic geometry approach quite adversely.
We emphasise that the main feature of Theorem 1.1 is that $\gamma_{n}>1/2$ since with $\gamma_{n}=n^{2}/(2n^{2}+2)<1/2$ it can be derived via the aforementioned elementary argument, avoiding the use of exponential sums.
@@
$\varPhi_{n}(H)=\sum_{\begin{subarray}{c}A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)\\ \det A\neq 0\end{subarray}}\frac{\varphi\left(|\det A|\right)}{|\det A|},$
where $\varphi(k)$ is the Euler function. Our results certainly apply to sums of many other multiplicative functions.
@@
$\varPhi_{n}(H)=2^{n^{2}}\sigma_{n}H^{n^{2}}+O\left(H^{n^{2}-\vartheta_{n}+o(1)}\right),$
@@
$\sigma_{n}=\prod_{p\penalty 10000\ \text{prime}}\,\left(1-\frac{1}{p}\right)\left(1+\frac{1}{p}\prod_{j=2}^{n}\left(1-p^{-j}\right)\right)$
@@
Our proof of Theorem 1.3 takes advantage of stronger bounds of exponential sums with square-free denominators with so-called monomial linear forms, see Lemma 3.7 below. We have not been able to establish a bound of similar strength in the case of denominators, which are squares of square-free integers, see, however, Lemma 3.8.
@@ 1.3. Notation and conventions
For a finite set $\mathcal{S}$} we use $\mathcal{S}$} to denote its cardinality.
@@
as $|k|\to\infty$, see [7, Equation (1.81)].
We also use the convention that if $A$ is an $n\times n$ matrix defined over ${\mathbb{Z}}_{m}$ or ${\mathbb{F}}_{p}$, then by $\det A=0$ we mean $\det A\equiv 0\bmod m$ or $\det A\equiv 0\bmod p$, respectively, which should be clear from the context. When $A$ has entries in an interval [H,H][-H,H], then any condition on the determinant modulo a positive integer $m$ or prime $p$ will be clearly indicated.
@@ 2.2. Matrices with fixed determinants
We need a bound on the number of matrices $A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right)$ with prescribed value of the determinant $\det A=d$. We recall that Duke, Rudnick and Sarnak [4], if $d\neq 0$, and Katznelson [9], when $d=0$, have obtained asymptotic formulas (with the main terms of orders $H^{n^{2}-n}$ and $H^{n^{2}-n}\log H$, respectively) for the number of such matrices when $d$ is fixed. However this is too restrictive for our purpose and so we use a uniform with respect to $d$ upper bound which is a special case of [15, Theorem 4].
@@ 3.1. Set-up
We are only interested in the sums $S_{m}(L)$ when the modulus $m=d$ or $m=d^{2}$ for a square-free integer $d\geq 1$. Thus we start with reduction to sums modulo $p$ or $p^{2}$. Namely, by the Chinese Remainder Theorem, we have the following decomposition, see [7, Equation (12.21)].
@@ 3.2. Linear sections of the variety of singular matrices
with at least one non-zero coefficient, say $a_{k,m}\neq 0$. Considering matrices with zero-entries everywhere, except at the $(k,m)$ position, which are clearly singular, we see that $a_{k,m}x_{k,m}=a_{0}$ for all $x_{k,m}\in K$, which is clearly impossible for $a_{k,m}\neq 0$.   $\sqcap\sqcup$
@@
Now, let $\mathbf{z}\neq\mathbf{0}$. Without loss of generality, we can assume that $\mathbf{z}=(z_{1},z_{2},\ldots,z_{n})$ with $z_{1}\neq 0$. Using elementary row operations, we can reduce the equation $\det\left(\mathbf{z}\mid\mathbf{Z}\right)=0$ to an equation $\det\left(\mathbf{z}_{0}\mid\mathbf{U}\right)=0$ for some matrix $\mathbf{U}$ and with $\mathbf{z}_{0}=(z_{1},0,\ldots,0)$. It is now obvious that $\det\left(\mathbf{z}_{0}\mid\mathbf{U}\right)=0$ holds for $p^{n-1}N_{n-1}(p)$ matrices $\mathbf{U}\in{\mathbb{F}}_{p}^{n\times(n-1)}$ with an arbitrary top row and and arbitrary singular $(n-1)\times(n-1)$ matrix over ${\mathbb{F}}_{p}$ on the other positions. Observe that row operations are invertible, and so for a given $\mathbf{z}$ each matrix $\mathbf{U}$ corresponds to a unique matrix $\mathbf{Z}$. Since there are $p^{n-1}-1$ choices for $\mathbf{z}\neq\mathbf{0}$ with $\ell(\mathbf{z})=0$, there are $(p^{n-1}-1)p^{n-1}N_{n-1}(p)$ such solutions.   $\sqcap\sqcup$
@@ 3.3. Bounding $S_{p}(L)$
Combining Lemma 3.2 with bounds of exponential sums along algebraic varieties, see, for example, [5, Proposition 1.2], one can immediately derive
@@
which after substitution in (3.3) concludes the proof.   $\sqcap\sqcup$
@@
for some linear forms $\ell$ and $L^{*}$ in $n$ and n(n1)n(n-1) variables, respectively, where, by our assumption, $\ell$ is nontrivial modulo $p$.
@@ 3.4. Bounding $S_{p^{2}}(L)$
for some linear forms $\ell$ and $L^{*}$ in $n$ and n(n1)n(n-1) variables, respectively, where by our assumption, $\ell$ is nontrivial modulo $p$.
@@
and substituting this bound in (3.10), we derive the desired result.   $\sqcap\sqcup$
@@ 3.5. Bounding $S_{d}(L)$ and $S_{d^{2}}(L)$
Substituting the estimates (3.12), (3.13) and (3.14) in (3.11), we derive the result.   $\sqcap\sqcup$
@@
Substituting the estimates (3.16), (3.17) and (3.18) in (3.15), we derive the result.   $\sqcap\sqcup$
@@ 4.1. Discrepancy and the Koksma-Szüsz inequality
We recall that the discrepancy $D(\Gamma)$ of a sequence $\Gamma=(\bm{\gamma}_{k})_{k=1}^{K}$ of $K$ points in the half-open $s$-dimensional unit cube $0,1)^{s}$, is defined by
@@
One of the basic tools in the uniformity of distribution theory is the celebrated Koksma–Szüsz inequality [10, 16] (see also [3, Theorem 1.21]), which links the discrepancy of a sequence of points to certain exponential sums.
@@ 4.2. Matrices with determinants with a divisibility condition in boxes
$N_{n}(m;H)=\#\{A\in{\mathcal{M}}_{n}\left({\mathbb{Z}};H\right):\penalty 10000\ \det A\equiv 0\bmod m\}.$
@@
$N_{n}(d;H)=N_{n}(d)\frac{(2H+1)^{n^{2}}}{d^{n^{2}}}+O\left(d^{n^{2}-2+1/n+o(1)}\right).$
@@
Let $m\geq 1$ be an integer and let $\mathbf{H}=\left(H_{ij}\right)_{i,j=1}^{n}$ with some positive integers $H_{ij}<m/2$ (below we set $m=d$). We extend the definition of $N_{n}(m;H)$ to $N_{n}(m;\mathbf{H})$ which is the number of $n\times n$ matrices with
$A=\left(a_{ij}\right)_{i,j=1}^{n}\in{\mathcal{M}}_{n}\left({\mathbb{Z}}\right),\qquad|a_{ij}|\leq H_{ij},\ i,j=1,\dots,n,$
@@
(4.1) $N_{n}(d;\mathbf{H})=N_{n}(d)\frac{1}{d^{n^{2}}}\prod_{i,j=1}^{n}\left(2H_{ij}+1\right)+O\left(d^{n^{2}-2+1/n+o(1)}\right).$
@@
For a positive integer $M\geq 1$, it is convenient to define ${\mathcal{L}}(M)$ as the set of all nonzero linear forms $L(\mathbf{X})\in{\mathbb{Z}}[\mathbf{X}]$ of the form (3.1) and with coefficients in the interval [M,M][-M,M].
@@
Hence, splitting each interval $0,H_{ij})$ into intervals of length $d$, and applying Lemma 4.1 to count points in incomplete boxes on the boundary, we have
$N_{n}(d;\mathbf{H})=N_{n}(d)\frac{1}{d^{n^{2}}}\prod_{i,j=1}^{n}\left(2H_{ij}+1\right)+O\left(M^{-1}N_{n}(d)+R\right),$
@@
(4.2) $N_{n}(d)\leq d^{n^{2}}\prod_{p\mid d}\left(p^{-1}+O\left(p^{-2}\right)\right)=d^{n^{2}-1+o(1)}.$
@@
(4.3) $N_{n}(d;\mathbf{H})=N_{n}(d)\frac{1}{d^{n^{2}}}\prod_{i,j=1}^{n}\left(2H_{ij}+1\right)+O\left(d^{n^{2}-1+o(1)}M^{-1}+R\right).$
@@
$\sum_{\begin{subarray}{c}L\in{\mathcal{L}}_{d}(M)\\ L\penalty 10000\ \text{monomial}\\ \gcd(L,d)=D\end{subarray}}\frac{1}{\rho(L)}\ll D^{-1}\log d.$
@@
(4.5) $R_{\text{mon}}=d^{n^{2}-n+o(1)}\sum_{\begin{subarray}{c}D\mid d\\ D\leq M\end{subarray}}D^{n}\sum_{\begin{subarray}{c}L\in{\mathcal{L}}_{d}(M)\\ L\penalty 10000\ \text{monomial}\\ \gcd(L,d)=D\end{subarray}}\frac{1}{\rho(L)}\leq d^{n^{2}-n+o(1)}M^{n-1}.$
@@
$\sum_{\begin{subarray}{c}L\in{\mathcal{L}}_{d}(M)\\ L\penalty 10000\ \text{non-monomial}\\ \gcd(L,d)=D\end{subarray}}\frac{1}{\rho(L)}\ll D^{-2}(\log d)^{n^{2}}.$
@@
(4.6) $\begin{split}R_{\text{non-mon}}&=d^{n^{2}-(n+1)/2+o(1)}\sum_{\begin{subarray}{c}D\mid d\\ D\leq M\end{subarray}}D^{(n+1)/2}\sum_{\begin{subarray}{c}L\in{\mathcal{L}}_{d}(M)\\ L\penalty 10000\ \text{non-monomial}\\ \gcd(L,d)=D\end{subarray}}\frac{1}{\rho(L)}\\ &\leq d^{n^{2}-(n+1)/2+o(1)}\left(M^{(n-3)/2}+1\right).\end{split}$
@@
$N_{n}(d;\mathbf{H})=N_{n}(d)\frac{1}{d^{n^{2}}}\prod_{i,j=1}^{n}\left(2H_{ij}+1\right)+O\left(d^{n^{2}+o(1)}\mathscr{E}\right),$
@@
Hence we now derive (4.1) and thus conclude the proof.   $\sqcap\sqcup$
@@
$N_{n}(d^{2};H)=N_{n}(d^{2})\frac{(2H+1)^{n^{2}}}{d^{2n^{2}}}+O\left(d^{2n^{2}-4(n+1)/(n+3)+o(1)}\right).$
@@
We follow the same path as in the proof of Lemma 4.2. In particular, we continue to use the notations $\mathbf{H}=\left(H_{ij}\right)_{i,j=1}^{n}$ and $N_{n}(m;\mathbf{H})$ (with $m=d^{2}$) and observe that it is enough to show that
(4.7) $\begin{split}N_{n}(d^{2};\mathbf{H})=N_{n}(d^{2})\frac{1}{d^{2n^{2}}}\prod_{i,j=1}^{n}&\left(2H_{ij}+1\right)\\ &+O\left(d^{2n^{2}-4(n+1)/(n+3)+o(1)}\right)\end{split}$
@@
$N_{n}(d^{2};\mathbf{H})=N_{n}(d^{2})\frac{1}{d^{2n^{2}}}\prod_{i,j=1}^{n}\left(2H_{ij}+1\right)+O\left(M^{-1}N_{n}(d^{2})+R\right),$
@@
(4.9) $N_{n}(d^{2};\mathbf{H})=N_{n}(d^{2})\frac{1}{d^{2n^{2}}}\prod_{i,j=1}^{n}\left(2H_{ij}+1\right)+O\left(M^{-1}d^{2n^{2}-2+o(1)}+R\right).$
@@
$N_{n}(d^{2};\mathbf{H})$ $=N_{n}(d^{2})\frac{1}{d^{2n^{2}}}\prod_{i,j=1}^{n}\left(2H_{ij}+1\right)$
@@
Choosing $M=d^{2(n-1)/(n+3)}$, we derive (4.7) and thus conclude the proof.   $\sqcap\sqcup$